OpenStudy (anonymous):

Find the domain of each function g(x)=5/x^2-25 ok I really don't remember how to do this, can someone please break it down for me quick and simple

OpenStudy (anonymous):

\[g(x)=5/( x^2-25)\]

OpenStudy (luigi0210):

set the denominator equal to 0. x^2-25=0 (x-5)(x+5)=0 ------------------------------ x-5=0 x+5=0 ------------------------------ x=?

OpenStudy (anonymous):

-5? @Luigi0210

OpenStudy (luigi0210):

..and?

OpenStudy (anonymous):

5

OpenStudy (luigi0210):

There you go

OpenStudy (luigi0210):

-5, and 5.

OpenStudy (anonymous):

ty

OpenStudy (luigi0210):

Yup :P

OpenStudy (anonymous):

@Luigi0210 what about on more complicated problems like f(t)=3t^2+5t+2

OpenStudy (anonymous):

then how could I find the domain again?

jimthompson5910 (jim_thompson5910):

Keep in mind that the values of 5 and -5 are not part of the domain since they make the denominator zero basically any number but 5 or -5 is in the domain

OpenStudy (luigi0210):

It would be all real numbers since there are no restrictions

jimthompson5910 (jim_thompson5910):

this is for the previous problem of 5/(x^2-25)

OpenStudy (anonymous):

ok so for the prvious problem domain=-5,5 would be wrong, I thought it was right

jimthompson5910 (jim_thompson5910):

quite the opposite actually

jimthompson5910 (jim_thompson5910):

the domain is anything but those two numbers

jimthompson5910 (jim_thompson5910):

I think you misread Luigi0210's answer a bit

OpenStudy (anonymous):

ok and what about the more complicated one?

jimthompson5910 (jim_thompson5910):

the domain is just the set of numbers that can be plugged into a function (where you will get some result out of it)

OpenStudy (anonymous):

that one would be all real numbers

jimthompson5910 (jim_thompson5910):

since plugging in x = 5 yields a division by zero error, it is NOT in the domain of f(x) = 5/(x^2-25)

jimthompson5910 (jim_thompson5910):

same thing with x = -5

OpenStudy (luigi0210):